Suppose that the steel gas tank in your car is completely filled when the temperature is 16 °C. How many gallons will spill out of the 88-Liters tank when the temperature rises to 40 °C?


1
Expert's answer
2021-11-14T17:11:18-0500

"ΔT = 40 -16 = 24 \; °C"

Change in volume of steel tank "= \u0394V_x"

Change in volume of gasoline "= \u0394V_g"

Coefficient of volume expansion of steel tank

"β_x = 30 \times 10^{-6} \; 1/°C"

Coefficient of volume expansion of gasoline

"β_g = 950 \times 10^{-6} \; 1/°C"

"V_0 = 88 \;L = 23.24 \;gal"

Change in volume of steel tank is

"\u0394V_x = \u03b2_xV_0\u0394T"

Change in volume of gasoline is

"\u0394V_g = \u03b2_gV_0\u0394T"

Required volume:

"ΔV = ΔV_g-ΔV_x \\

ΔV=(β_g -β_x)V_0ΔT \\

ΔV=(950 \times 10^{-6} -36 \times 10^{-6}) (23.24 \;gal) (24 \; °C) \\

ΔV= 0.509 \;gal"

Answer: 0.509 gal